You are tiling a bathroom floor with tiles that are regular hexagons.
Each tile has 6 inch sides.
How many tiles will you need if the floor of the bathroom is 100 square feet?

Answer by mananth(16946)

The area of ar regular hexan is given by the formula 3sqrt3 / 2 * s^2 . s= side
convert the side to feet = 1/2 feet
=3* sqrt3 * (1/2)^2 / 2
= 0.649 square feet
.
Area of floor = 100
Area of each tile = 0.549 sq.feet
No of tiles = 100 / 0.649
= 154 tiles